A radiator contains 15 quarts of fluid, 15% of which is antifreeze. How much fluid should be drained and replaced with pure antifreeze so that the new mixture is 50% antifreeze? (Round your answer to two decimal places.)
Answer by richwmiller(17219)

0.15*15-0.15x+1x=0.5(15)
2.25+0.85x=7.5
0.85x=7.5-2.25
0.85x=5.25
x=6.17647059
x=6.18 quarts rounded to two decimal places
check
0.15*8.82352941+1*6.17647059=0.5(15)
1.32352941+6.17647059=7.5
7.5=7.5
ok
coderep